solve the system using substitution




{5+10x=y
{20+5x=y

Answer by stanbon(48573)

solve the system using substitution
---------------
{5+10x=y
{20+5x=y
-------
5 + 10x = 20+5x
5x = 15
x = 3
=============
Cheers,
Stan H.